Literaturverzeichnis 149

[30] A. M. Ozorio de Almeida, Hamiltonian Systems: Chaos and Quantization (Cambridge University Press, Cambridge, 1988).

[31] K. R. Meyer und G. R. Hall, Introduction to Hamiltonian Dynamical Systems and the N-Body Problem, Band 90 der Reihe Applied Mathematical Sciences (Springer, New York, 1992).

[32] A. J. Lichtenberg und M. A. Lieberman, Regular and Chaotic Dynamics, Band 38 der Reihe Applied Mathematical Sciences (Springer, New York, 1992).

[33] K. R. Meyer, Trans. Am. Math. Soc. 149, 95 (1970).

[34] A. D. Bruno, Math. USSR Sbornik 12, 271 (1970).

[35] A. D. Bruno, preprint Nr. 18, Inst. Prikl. Mat. Akad. Nauk SSSR, Moskau (in Russisch) (1972).

[36] A. M. Ozorio de Almeida und J. H. Hannay, J. Phys. A 20, 5873 (1987).

[37] M. Sieber, J. Phys. A 29, 4715 (1996).

[38] D. A. Sadovskií, J. A. Shaw und J. B. Delos, Phys. Rev. Lett. 75, 2120 (1995).

[39] D. A. Sadovskií und J. B. Delos, Phys. Rev. E 54, 2033 (1996).

[40] E. Bogomolny, O. Bohigas und P. Leboeuf, Phys. Rev. Lett. 68, 2726 (1992).

[41] E. T. Copson, Asymptotic Expansions (Cambridge University Press, Cambridge, 1965).

[42] M. V. Fedoryuk, in Analysis I, Band 13 der Reihe Encyclopaedia of Mathematical Sciences, herausgegeben von R. V. Gamkrelidze (Springer, Berlin Heidelberg, 1989).

[43] R. Balian und C. Bloch, Ann. Phys. (N. Y.) 85, 514 (1974).

[44] M. V. Berry, Proc. R. Soc. Lond. A 422, 7 (1989).

[45] P. A. Boasman und J. P. Keating, Proc. R. Soc. Lond. A 449, 629 (1995).

[46] R. Thom, Structural Stability and Morphogenesis (Benjamin, Reading, Mass., 1975).

[47] J. J. Duistermaat, Commun. Pure Appl. Math. 27, 207 (1974).

[48] J. N. L. Connor, Molec. Phys. 31, 33 (1976).

[49] M. V. Berry, Adv. Phys. 25, 1 (1976).

[50] T. Poston und I. N. Stewart, Catastrophe Theory and its Applications (Pitman, London, 1978).

[51] V. I. Arnol’d, in Dynamical Systems V, Band 5 der Reihe Encyclopaedia of Mathematical Sciences, herausgegeben von V. I. Arnol’d (Springer, Berlin Heidelberg, 1994).

[52] F. J. Dyson, J. Math. Phys. 3, 140 (1962).

[53] M. L. Mehta, Random Matrices (Academic Press, New York, 1991).

[54] O. Bohigas, M. J. Giannoni und C. Schmit, Phys. Rev. Lett. 52, 1 (1984).

[55] F. Haake, M. Kuś, H.-J. Sommers, H. Schomerus und K. Życzkowski, J. Phys. A 29, 3641 (1996).

[56] L. G. Yaffe, Rev. Mod. Phys. 54, 407 (1982).